OpenStudy (anonymous):

Solve –2(2x – 3) = –6

OpenStudy (anonymous):

x=3

OpenStudy (angela210793):

-4x+6=-6 -4x=-12 x=3

OpenStudy (aroub):

-4x+6=-6 -4x=-6-6 -4x=-12 x=3

OpenStudy (anonymous):

x=3